OpenStudy (anonymous):

Using the ratio of perfect squares method , what is √14 rounded to the nearest hundredth? Will FAN and MEDAL!

OpenStudy (anonymous):

I don't think I've ever heard of this method

OpenStudy (anonymous):

I can help you solve it, but I am not sure if I will be doing it by this method

OpenStudy (anonymous):

umm can you tell us what the perfect square method is?

OpenStudy (anonymous):

A number made by squaring a whole number. 16 is a perfect square because 42 = 16 25 is also a perfect square because 52 = 25 etc

OpenStudy (anonymous):

Is it when you find like the closest perfect squares to the non-perfect square you are working with

OpenStudy (anonymous):

Like say we have to do the same thing, but with \(\sqrt{23}\) The closest perfect squares to that is \(\sqrt{16}\) and \(\sqrt{25}\) So \(\sqrt{16}\) < \(\sqrt{23}\) < \(\sqrt{25}\) Now you simplify the two perfect squares. 4 < \(\sqrt{23}\) < 5

OpenStudy (anonymous):

I am just going to try to put in the square root of 14 and round it to the nearest hundreth which would be 3.74